OpenStudy (anonymous):

A map is scaled so that 2 cm on the map is equal to 10 actual miles. If two cities on the map are 20 cm apart, proportion would you use to solve the problem? A. the fraction of map over actual equals twenty-tenths equals the fraction x over 2 B. the fraction of map over actual equals the fraction 2 over x equals the fraction twenty-tenths C. the fraction of map over actual equals two-tenths equals the fraction 20 over x D. the fraction of map over actual equals the fraction x over 10 equals twenty-halves

OpenStudy (anonymous):

A. map/actual = 20/10 = x/2 B. map/actual = 2/x = 20/10 C. map/actual = 2/10 = 20/x D. map/actual = x/10 = 20/2

hartnn (hartnn):

since map=2cm and actual = 10 miles map / actual = 2/10 ok? now 2o cm in map , so 20 would come in numerator x in actual ,so x comes in denominator.... so your proportion is 2/10 = 20/ x does this make sense ?

hartnn (hartnn):

which is option C.